OpenStudy (anonymous):

if x^2+mx+m is a perfect square trinomial, which equation must be true

OpenStudy (anonymous):

where are the choices?

OpenStudy (mathmate):

hint: \[(a+b)^2=a^2+2ab+b^2\]

OpenStudy (anonymous):

One way to do this is to complete the square: \[\begin{align*}x^2+mx+m&=x^2+mx+\frac{m^2}{4}-\frac{m^2}{4}+m\\ &=\left(x+\frac{m}{2}\right)^2-\frac{m^2}{4}+m\\ &=\left(x+\frac{m}{2}\right)^2+\frac{4m-m^2}{4} \end{align*}\] Since this is a perfect square trinomial, you must have \[\frac{4m-m^2}{4}=0\] Otherwise, you wouldn't be able to write \(x^2+mx+m\) as a squared binomial.